OpenStudy (anonymous):

y=-2x+65 is??? someone help please!

OpenStudy (smileyxl3):

What are you trying to find? what the values of x and y are?

OpenStudy (anonymous):

this is the full question. The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later.

OpenStudy (anonymous):

@smileyxl3

OpenStudy (smileyxl3):

y=65-2x put x = 6 y = 65 - 2(6) = 65 -12 =53 degrees

OpenStudy (anonymous):

Okay thank you (: